OpenStudy (dtan5457):

2x+4/radical x+2 Which is equivalent to this? Does this equal to 2 radical x+2? If so, can someone demonstrate why?

OpenStudy (anonymous):

\[\frac{2x+4}{\sqrt{x+2}}\]?

OpenStudy (dtan5457):

yes

OpenStudy (anonymous):

\[\frac{2(x+2)}{\sqrt{x+2}}\]

OpenStudy (anonymous):

is a start

OpenStudy (anonymous):

and it is always the case that \[\frac{z}{\sqrt{z}}=\sqrt z\] you can check that easily by 1) squaring both sides and see that you get \(z\) or 2) write in exponential notation and subtract the exponents

OpenStudy (anonymous):

so yes, \[\frac{2(x+2)}{\sqrt{x+2}}=2\sqrt{x+2}\]

OpenStudy (dtan5457):

is that a specific formula?

OpenStudy (anonymous):

no it is not a formula, it is just a fact

OpenStudy (anonymous):

\[\frac{9}{\sqrt9}=\frac{9}{3}=3=\sqrt9\] for example

OpenStudy (dtan5457):

Alright I get that now. Mind if you help me with an inverse problem real quick? No one answered me correctly when I asked it.

OpenStudy (anonymous):

\[\frac{x}{\sqrt{x}}=\frac{x}{x^{\frac{1}{2}}}=x^{1-\frac{1}{2}}=x^{\frac{1}{2}}=\sqrt{x}\]

OpenStudy (dtan5457):

y=3x+2/4x-7

OpenStudy (anonymous):

i can try

OpenStudy (anonymous):

LOL no wonder no one wanted to do it, is a bunch of pain in the butt algebra solve \[x=\frac{3y+2}{4y-7}\] for \(y\) takes a bunch of steps

OpenStudy (anonymous):

we can walk through them if you like

OpenStudy (dtan5457):

i would love to

OpenStudy (anonymous):

1) multiply both sides by \(4y-7\) to clear the denominator let me know what you get

OpenStudy (dtan5457):

yeah thats what they told me before but i dont necessarily get why it clears the denominator i thought the conjugate does

OpenStudy (anonymous):

you are not rationalizing a denominator you are solving if you had \[\frac{3y+2}{4y-7}=5\] and had to solve for \(y\) what would you do?

OpenStudy (dtan5457):

make the fraction a whole number

OpenStudy (anonymous):

what do you mean by that? can you solve this \[\frac{3y+2}{4y-7}=5\] ?

OpenStudy (anonymous):

how about this \[\frac{2x}{3}=7\] what would you do here?

OpenStudy (dtan5457):

i dont know if anything i almost have no clue in is taking out denominators

OpenStudy (anonymous):

ok how about this \[\frac{x}{3}=5\]

OpenStudy (anonymous):

multiply 5 by 3 and get \[x=5\times 3=15\]

OpenStudy (anonymous):

\[\frac{2x+1}{3}=5\\ 2x+1=3\times 5=15\\ 2x=14\\ x=7\]

OpenStudy (dtan5457):

oh....

OpenStudy (anonymous):

first step is always to multiply by the denominator so you can solve the equation \[\frac{x+1}{x-3}=7\\ x+1=7(x-3)\\ x+1=7x-21\\ etc\]

OpenStudy (dtan5457):

ok i get that now

OpenStudy (anonymous):

want to try this one \[\frac{3y+2}{4y-7}=5\]

OpenStudy (dtan5457):

3y+2=20y-35 -33=17y

OpenStudy (dtan5457):

i mean 37=17y

OpenStudy (anonymous):

ok now we get to the real question \[\frac{3y+2}{4y-7}=x\] which is harder because we have x and not 5, but it is the same steps exactly multiply by \(4y-7\) to get it all on one line

OpenStudy (anonymous):

what do you get?

OpenStudy (dtan5457):

3y+2=4xy-7x

OpenStudy (dtan5457):

shouldn't you switch the variables first?

OpenStudy (anonymous):

ok good now we are solving for \(y\) so we need all terms with \(y\) on one side of the equal sign, everything else on the other yes, we switched already

OpenStudy (anonymous):

\[3y+2=4xy-7x\] your choice as to where you want the y terms, right or left, makes no difference pick one

OpenStudy (dtan5457):

how do you minus the 4xy part again?

OpenStudy (anonymous):

you want them on the left i guess you just write it

OpenStudy (anonymous):

\[3y+2=4xy-7x\\ 3y+2-4xy=-7x\]

OpenStudy (anonymous):

we got three steps to go since you put the y terms on the left, put everything without a y in it on the right

OpenStudy (dtan5457):

well thats done right

OpenStudy (anonymous):

uh no

OpenStudy (anonymous):

you have a 2 on the left still

OpenStudy (dtan5457):

couldn't you just divide everything by -7 to get x now

OpenStudy (anonymous):

we are solving for y no you do not divide until the very last step

OpenStudy (dtan5457):

ok so 3y-4xy=-7x-2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now we are solving for \(y\) and you have the two terms with \(y\) on one side of the equal sign factor it out

OpenStudy (anonymous):

you know what i mean by that?

OpenStudy (dtan5457):

yea but i dont know how to do it

OpenStudy (anonymous):

\[3y-4xy=-7x-2\\ (3-4x)y=-7x-2\]

OpenStudy (anonymous):

and now the LAST step is to divide to get \(y\) \[y=\frac{-7x-2}{3-4x}\]

OpenStudy (anonymous):

which i would probably write as \[y=\frac{7x+2}{4x-3}\]

OpenStudy (dtan5457):

i think i just died a little

OpenStudy (anonymous):

you want to do it again?

OpenStudy (anonymous):

or we could do this one \[\frac{3y+2}{4y-7}=5\] again and see that the steps were identical

OpenStudy (dtan5457):

i think the main problem is how to cancel the denominator. if you look at this other problem i asked http://openstudy.com/users/dtan5457#/updates/547d3062e4b0ae92adfff782 why did they mutliply by the conjugate, and the denominator goes away in that situation?

OpenStudy (dtan5457):

OK THANKS I ACTUALLY GOT IT NOW